OpenStudy (anonymous):

An electron with speed 2.25×107m/s is traveling parallel to a uniform electric field of magnitude 1.10×104N/C .? A) How far will the electron travel before it stops? B) How much time will elapse before it returns to its starting point?

OpenStudy (aaronq):

\(F=qE=ma\) then use kinematic equations

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

I keep getting 2.4 x 10^-6 and it's wrong

OpenStudy (aaronq):

the acceleration is \(a=\dfrac{qE}{m}=\dfrac{1.6*10^{-19}~C*1.10×10^4~N/C}{9.1*10^{-31}~kg}=1.93*10^{15}~m/s^2\) Use this acceleration in the kinematic equation of motion under constant acceleration. NOTE: it's negative because of the direction of the electron.

OpenStudy (anonymous):

what is the kinematic equation of motion under constant acceleration?

OpenStudy (aaronq):

Kinematic equations \(\sf v_f=v_i+a\Delta t\) \(\sf x_f=x_i+v_o\Delta t +\dfrac{1}{2}a_x(\Delta t)^2\) \(\sf v_f^2=v_i^2+2a_x\Delta x\)

OpenStudy (anonymous):

so, what values would you use for delta x and for vf and vi

OpenStudy (aaronq):

Have you done questions in mechanics where it asks "how high does an object go when you throw it upwards with a velocity of x m/s"? It's the same thing.

OpenStudy (aaronq):

I'll help you with the first one. Use: \(\sf v_f^2=v_i^2+2a_x\Delta x\) \(\sf(0)^2=( 2.25*10^7~m/s )^2+2(-1.93∗10^{15} m/s^2)\Delta x\) \(\sf \Delta x =0.1311528~m \approx 0.131 ~m\)

OpenStudy (anonymous):

thank you so much! now to find out how much time is elapsed, do you just use vy=at?

OpenStudy (aaronq):

not quite. It's asking for the time of the ROUND trip (there and back). The accelerations are in different directions. First negative then positive 

OpenStudy (aaronq):

wait i got the electric field the opposite way, but it makes no difference.

OpenStudy (anonymous):

I'm sorry, I'm confused so what number do you use for your v and will it be a positive number or a negative number?

OpenStudy (anonymous):

i understand the two equations a=ve and a=-ve, but what values do you use for v

OpenStudy (aaronq):

okay, so we got the distance 0.131 m, the initial speed 2.25×10^7m/s, and the acceleration 1.93*10^15 m/s^2. This is parabolic motion,  \(\sf x_f=x_i+v_o\Delta t +\dfrac{1}{2}a_x(\Delta t)^2\) \(\sf 0 =(2.25×10^7m/s)\Delta t -\dfrac{1}{2}(1.93*10^{15} m/s^2)(\Delta t)^2\) \(\sf \Delta t=0.0000000233160622~s=2.33*10^{-8}~s=23.2 ~ns\)